OpenStudy (anonymous):

Solve the SYSTEM!!!!! 2x - 4y = 4 x + 4y = 14 a. 2,6 b. 4,6 c. 6,2 d. 6,4

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (texaschic101):

substitution is probably easier in this problem...so here we go.. x + 4y = 14 --- subtract 4y from both sides x = -4y + 14 I wanted to isolate a variable and x was the easiest to isolate. So now we know x = -4y + 14....so sub that in for x in the other equation. 2x - 4y = 4 2(-4y + 14) - 4y = 4 --- distribute through the parenthesis -8y + 28 - 4y = 4 -- subtract 28 from both sides -8y - 4y = 4 - 28 -- combine like terms -12y = -24 -- divide both sides by -12 y = 2 now sub 2 in for y in either equation to find x x + 4y = 14 x + 4(2) = 14 x + 8 = 14 x = 14 - 8 x = 6 so your solution is : (6,2)

OpenStudy (anonymous):

thanks so much!!!!!